# ScholarGym: Benchmarking Deep Research Workflows on Academic Literature Retrieval

### v1

@@ Abstract
Tool-augmented large language models have advanced from single-turn question answering to deep research workflows that iteratively plan queries, invoke external tools, and synthesize information to address complex information needs. Evaluating such workflows presents a fundamental challenge: reliance on live APIs introduces non-determinism, as tool invocations may yield different results across runs due to temporal drift, rate limiting, and evolving backend states. This variance undermines reproducibility and invalidates cross-system comparisons.
We present ScholarGym, a simulation environment for reproducible evaluation of deep research workflows on academic literature. The environment decouples workflow components into query planning, tool invocation, and relevance assessment, enabling fine-grained analysis of each stage under controlled conditions. Built on a static corpus of 570K papers with deterministic retrieval, ScholarGym provides 2,536 queries with expert-annotated ground truth. Experiments across diverse backbone models reveal how reasoning capabilities, planning strategies, and selection mechanisms interact over iterative refinement.
@@ 1 Introduction
Large language models (LLMs) have evolved from single-turn question answering to sophisticated workflows that plan queries, invoke external tools, and synthesize retrieved information (Nakano et al., 2021; Schick et al., 2023). When applied to academic literature, these deep research workflows decompose complex research questions into subqueries, retrieve candidate papers through multiple tool calls, and filter results based on relevance criteria. Commercial systems such as OpenAI Deep Research and Gemini Deep Research (OpenAI, 2025a; Google, 2025a) have demonstrated the practical value of this paradigm.
Evaluating deep research workflows, however, presents a fundamental challenge. Existing approaches rely on live APIs that introduce non-determinism: tool invocations may yield different results across runs due to temporal drift in search indices, rate limiting that causes incomplete responses, and evolving backend states that alter ranking behavior. This variance undermines reproducibility. Reported performance cannot be verified independently, cross-system comparisons become invalid, and measuring progress in the field is difficult.
To address these issues, we introduce ScholarGym, a simulation environment for reproducible evaluation of deep research workflows on academic literature. By replacing live APIs with a static corpus and deterministic retrieval, ScholarGym isolates workflow capabilities from environmental noise. The environment decouples the workflow into three stages—query planning, tool invocation, and relevance assessment—enabling fine-grained analysis of each component under controlled conditions.
Our contributions are as follows:
(1) An iterative deep research workflow. We formalize deep research as a modular process comprising query planning, tool invocation, and relevance assessment, underpinned by a memory mechanism that maintains reasoning state across iterations.
(2) A reproducible simulation environment. We introduce ScholarGym, which grounds this workflow in a static corpus of 570K papers with deterministic retrieval, decoupling algorithmic reasoning from live API variance to enable rigorous evaluation of long-horizon research systems.
(3) Systematic experiments. We evaluate state-of-the-art LLMs, revealing that “extended thinking” paradigms induce a distinct precision-recall trade-off. We further identify query planning and relevance assessment as dual bottlenecks limiting open-source performance on complex research inquiries.
@@ 2.1 Deep Research Evaluation
Benchmarks for deep research have evolved from factoid QA to multi-step information synthesis. GAIA (Mialon et al., 2023) and BrowseComp (Wei et al., 2025) evaluate tool use and web navigation but focus on intermediate reasoning rather than comprehensive outputs. DeepResearch Bench (Du et al., 2025; Li et al., 2026), Rigorous Bench (Yao et al., 2025), DEER (Han et al., 2025), and FINDER (Zhang et al., 2025a) shift toward holistic evaluation of report quality using multidimensional rubrics.
A shared limitation is reliance on live APIs. Chen et al. (2025) observe that temporal drift in web content creates a “reproducibility gap” obscuring algorithmic progress. Zhang et al. (2025a) report that evaluation failures frequently stem from anti-scraping mechanisms and URL timeouts rather than model deficiencies. LiveResearchBench (Wang et al., 2025) acknowledges that the absence of fixed ground truth renders verification costly and noisy. These observations motivate controlled evaluation that decouples workflow capabilities from environmental variance.
@@ 2.2 Academic Literature Retrieval
Static benchmarks offer a partial remedy. LitSearch (Ajith et al., 2024) constructs a fixed corpus from ACL and ICLR papers and finds that dense retrievers outperform commercial search engines on deep understanding queries. SciNetBench (Shao et al., 2025) uses an OpenAlex snapshot for relation-aware retrieval. HiSciBench (Zhang et al., 2025b) highlights that literature review generation suffers from factuality gaps exacerbated by dynamic sources. Our work extends this direction by coupling a static corpus with an iterative workflow, enabling reproducible assessment of multi-turn planning and selection.
@@ 2.3 Deep Research Systems
RL-based systems train agents end-to-end on search trajectories. DeepResearcher (Zheng et al., 2025) applies reinforcement learning in real-world environments. Search-R1 (Jin et al., 2025) and SimpleTIR (Xue et al., 2025) train on multi-turn tool interactions. DeepDive (Lu et al., 2025) integrates knowledge graphs with multi-turn RL. Gao et al. (2025) and Li et al. (2025) scale asynchronous RL for long-horizon search. Tongyi DeepResearch (Team et al., 2025) validates algorithms on a Wikipedia snapshot to address non-stationarity—a “wind tunnel” strategy that parallels our motivation. Step-DeepResearch (Hu et al., 2025a) synthesizes atomic capabilities to bypass unstable API feedback.

### v2

# ScholarGym: Benchmarking Large Language Model Capabilities in the Information-Gathering Stage of Deep Research

@@ Abstract
Large language models have advanced from single-turn question answering to deep research systems that iteratively decompose research questions, invoke retrieval tools, and synthesize information across multiple rounds. Evaluating such systems typically involves scoring their final research reports holistically, but this end-to-end paradigm tightly couples the language model’s decision-making, workflow design, and environmental feedback, precluding decomposable analysis of individual components.
[…]
@@ 1 Introduction
In recent years, large language models (LLMs) have evolved from single-turn question answering to deep research systems capable of long-horizon, iterative information gathering and synthesis (Li et al., 2025c; Wu et al., 2025; Zhang et al., 2025a). In academic literature retrieval, such systems progressively decompose a research question into actionable subqueries, repeatedly invoke retrieval tools, and dynamically adjust their search and filtering strategies through multi-round feedback, ultimately integrating relevant findings into structured research reports or survey-style analyses (Wang et al., 2024; Bao et al., 2025; He et al., 2025). Commercial systems operating in this paradigm—including OpenAI Deep Research, Gemini Deep Research, and Kimi Agent (OpenAI, 2025a; Google, 2025a; Kimi, 2025)—have demonstrated substantial practical value.
As deep research systems advance rapidly, numerous efforts have attempted to evaluate this class of tasks, typically by running a complete deep research pipeline and scoring the final research report holistically (Gou et al., 2025; Du et al., 2025; Li et al., 2025b). While such evaluations provide useful reference points for overall system performance under particular configurations, their principal limitation is that multiple critical factors—the language model’s decision-making, tool interface design, workflow orchestration, and environmental feedback—remain tightly coupled throughout evaluation. These factors mutually influence and constrain one another, making it difficult to attribute observed outcomes to specific components. Under this setup, evaluation results can only indicate whether the system works as a whole, but cannot answer which components are effective and which decisions truly determine research quality.
A principled evaluation framework for deep research should satisfy several key requirements. First, evaluation should center on the language model’s decisions and actions, enabling the contribution of each component to be explicitly disentangled and analyzed, rather than relying solely on end-to-end judgment. Second, since final reports vary widely in structure and style, evaluation should target the upstream behaviors that most determine report quality—information gathering, coverage, and filtering—that have a decisive impact on research quality. Third, to support reproducible and comparable analysis, the evaluation environment should eliminate the non-determinism introduced by live web APIs and open-ended search, characterizing model and component interactions under controlled conditions.
We introduce ScholarGym, an evaluation environment that targets the information-gathering stage of deep research on academic literature. ScholarGymfixes a unified workflow and decomposes it into three stages—Query Planning, Tool Invocation, and Relevance Assessment—with the language model as the sole decision-maker at each stage. All retrieval operates over a static corpus of 570K papers with deterministic ranking, enabling systematic comparison of model behavior during information gathering. We construct 2,536 queries with expert-annotated ground truth and equip the workflow with a memory mechanism that maintains reasoning state across iterations, enabling rigorous, reproducible evaluation of long-horizon academic retrieval.
[…]
@@ 2.1 Deep Research Evaluation
Benchmarks for deep research have evolved from factoid QA to multi-step information synthesis. GAIA (Mialon et al., 2023) and BrowseComp (Wei et al., 2025) evaluate tool use and web navigation but focus on intermediate reasoning rather than comprehensive outputs. DeepResearch Bench (Du et al., 2025; Li et al., 2026), Rigorous Bench (Yao et al., 2025), DEER (Han et al., 2025), and FINDER (Zhang et al., 2025b) shift toward holistic evaluation of report quality using multidimensional rubrics.
A shared limitation is reliance on live APIs. Chen et al. (2025) observe that temporal drift in web content creates a “reproducibility gap” obscuring algorithmic progress. Zhang et al. (2025b) report that evaluation failures frequently stem from anti-scraping mechanisms and URL timeouts rather than model deficiencies. LiveResearchBench (Wang et al., 2025) acknowledges that the absence of fixed ground truth renders verification costly and noisy. These observations motivate controlled evaluation that decouples workflow capabilities from environmental variance.
@@ 2.2 Academic Literature Retrieval
Static benchmarks offer a partial remedy. LitSearch (Ajith et al., 2024) constructs a fixed corpus from ACL and ICLR papers and finds that dense retrievers outperform commercial search engines on deep understanding queries. SciNetBench (Shao et al., 2025) uses an OpenAlex snapshot for relation-aware retrieval. HiSciBench (Zhang et al., 2025c) highlights that literature review generation suffers from factuality gaps exacerbated by dynamic sources. Our work extends this direction by coupling a static corpus with an iterative workflow, enabling reproducible assessment of multi-turn planning and selection.
@@ 2.3 Deep Research Systems
RL-based systems train agents end-to-end on search trajectories. DeepResearcher (Zheng et al., 2025) applies reinforcement learning in real-world environments. Search-R1 (Jin et al., 2025) and SimpleTIR (Xue et al., 2025) train on multi-turn tool interactions. DeepDive (Lu et al., 2025) integrates knowledge graphs with multi-turn RL. Gao et al. (2025) and Li et al. (2025a) scale asynchronous RL for long-horizon search. Tongyi DeepResearch (Team et al., 2025) validates algorithms on a Wikipedia snapshot to address non-stationarity—a “wind tunnel” strategy that parallels our motivation. Step-DeepResearch (Hu et al., 2025a) synthesizes atomic capabilities to bypass unstable API feedback.